# Bias in Language Models: Beyond Trick Tests and Toward RUTEd Evaluation

### v1

@@ Abstract
Bias benchmarks are a popular method for studying the negative impacts of bias in LLMs, yet there has been little empirical investigation of whether these benchmarks are actually indicative of how real world harm may manifest in the real world. In this work, we study the correspondence between such decontextualized “trick tests” and evaluations that are more grounded in Realistic Use and Tangible Effects (i.e. RUTEd evaluations). We explore this correlation in the context of gender-occupation bias–a popular genre of bias evaluation. We compare three de-contextualized evaluations adapted from the current literature to three analogous RUTEd evaluations applied to long-form content generation. We conduct each evaluation for seven instruction-tuned LLMs. For the RUTEd evaluations, we conduct repeated trials of three text generation tasks: children’s bedtime stories, user personas, and English language learning exercises. We found no correspondence between trick tests and RUTEd evaluations. Specifically, selecting the least biased model based on the de-contextualized results coincides with selecting the model with the best performance on RUTEd evaluations only as often as random chance. We conclude that evaluations that are not based in realistic use are likely insufficient to mitigate and assess bias and real-world harms.
large language models, algorithmic fairness, algorithmic bias, fairness, bias, Llama 2, Flan-PaLM 2, metrics, evaluations, benchmarks
[…]
@@ 1 Introduction
As LLMs are increasingly used in everyday life, concerns have naturally been raised about the impacts these models have on their users and society at large. From these concerns have sprung a number of benchmarks to assess bias in LLMs (for a recent survey, see Gallegos et al., 2023). However, many of the common benchmarks for evaluating bias in LLMs are not based on realistic use of the models and have little justification for how the bias they purport to measure would manifest as real-world harm. In their review of how the NLP literature uses the term “bias,” Blodgett et al. (2020) found that papers use a wide range of normative motivations–often only briefly or vaguely specified–including stereotyping, questionable correlations between model behavior and language features, allocational harms (e.g., the distribution of jobs or financial resources), and a nebulous category of other representational harms (e.g., system performance, misrepresentation, denigration). And yet, the benchmarks from this literature are a mainstay of bias assessment for LLMs deployed in the real world. For example, the PaLM 2 models developed by Google and the Claude models developed by Anthropic were both tested with the Bias Benchmark for QA (BBQ), including a remark from Anthropic on their improvement in BBQ performance from Claude 1 to Claude 2 (Anthropic, 2023; Google, 2022).
We have very little understanding of how well bias benchmarks correlate with real-world bias and harm, particularly in the contextually rich use case of text generation. In this paper, we fill this gap. We study whether typical bias benchmarks are useful for evaluating how and whether bias manifests in more realistic applications. In particular, we argue that most existing bias benchmarks are trick tests — de-contextualized evaluations based on contrived scenarios designed more to elicit the undesirable “biased” behavior than to serve as a genuine test of performance in a realistic setting for that task. We hold these tests in contrast to evaluations that are are grounded in Realistic Use and Tangible Effects, or RUTEd (“rooted”) evaluations. The need for RUTEd evaluations echoes recent calls for more sociotechnical evaluations of AI (Weidinger et al., 2023) as also articulated in a recent talk by Deb Raji (“Grounded Evaluations for Assessing Real-World Harms” at a recent NeurIPS workshop). We compare measurements of bias in trick tests to analogous measures of bias in RUTEd evaluations. We find that, if we use trick test evaluations to select which candidate model is the least biased, we would do no better than chance at selecting the model with the least bias in more RUTEd long-form content generation. Further, bias evaluations in each RUTEd use case are largely uncorrelated with each other, suggesting the need for bespoke evaluations of LLM bias, tailored to each highly specific context of use.
In this paper, we use the following terminology. An evaluation is composed of a metric as applied to the output of a model performing some task on some corpus of test cases. The task defines the model’s objective, and the corpus of test cases for that task are provided in a dataset. A metrics is a quantitative formulas that summarizes the model’s performance at that task for that dataset. When an evaluation becomes standardized, it is a benchmark. Evaluations, including benchmarks, can have varying degrees of connection to realistic use and harms.
2 The Evolution of Bias Evaluations
Bias evaluations are typically divided into two categories: intrinsic and extrinsic. Intrinsic bias evaluations measure properties inherent to the model in question. Extrinsic bias evaluations measure the model’s biases relative to a specified task. While this distinction was once clear, the advancing frontier of language modeling has added more stages of model development (e.g., fine-tuning, in-context learning), the bound of what can be considered “intrinsic” has expanded, such that task performance that for previous generations of language models would have been considered “intrinsic” is now sometimes considered an inherent property of the model.
2.1 Static word embedding models
@@
Two popular early examples of intrinsic bias metrics are the Word Embedding Association Test (WEAT) benchmark Caliskan et al. (2017)) and the similar approach of Bolukbasi et al. (2016). Both aggregate cosine similarity measures between words associated with different identity groups (e.g., “he,” “she”) with words in a domain of interest (e.g., occupations). In the paper that introduced the intrinsic-extrinsic dichotomy (Goldfarb-Tarrant et al., 2020), the intrinsic metric of WEAT (in both English and Spanish) was contrasted with extrinsic metrics based on models that used those embeddings for the tasks of coreference resolution and hate speech detection.
@@ 2.2 LLMs
As the dominant paradigm shifted towards LLMs, so did conceptualization of what is considered “intrinsic.” In contrast to static word embedding models, LLMs contain dynamic embeddings that change with context. By design, they are inherently capable of next-word prediction, a task that can be applied to numerous particular objectives. To evaluate bias in this paradigm, Guo & Caliskan (2021) developed an extension of WEAT, the Contextualized Embedding Association Test (CEAT). Another paper on the intrinsic-extrinsic connection, Cao et al. (2022), adapted to this shifting paradigm with numerous experiments on 19 models, primarily variants of BERT and GPT-2. In this study, they considered CEAT and two other benchmarks–StereoSet (Nadeem et al., 2020) and ILPS (Kurita et al., 2019)–as “intrinsic metrics,” even though they are not based on the embedding space itself but on the log probabilities of words in text that can evoke stereotypes. These probabilities constitute task performance in the sense that they reflect the next-word predictions of a non-zero temperature LLM over many trials.
Popular examples of other probability-based bias evaluations include CrowS-Pairs and several in the BIG-bench benchmarking suite. The CrowS-Pairs benchmark compares the probabilities of stereotype-related tokens conditional on the presence of identity-related tokens. For example, it compares the probabilities of the tokens People who live in <modified> are alcoholics when the modified part is trailer parks to the probabilities when it is mansions, referring to a stereotype of alcoholism and low socioeconomic status (Nangia et al., 2020). Token probabilities can also be calculated across sentences, as in the StereoSet benchmark, which presents the model with a sentence referring to a particular racial group in a way that may evoke stereotypes. Here,the metric is based on the probability that a particular stereotype-related sentence follows (Nadeem et al., 2020). Several subtasks in the popular BIG-bench benchmarking suite related to bias are based on similar probabilities (Srivastava et al., 2022). For example, the Gender Sensitivity English evaluation is based on the next-word probability in a simple sentence like Sailor is. The sum of the next-word probabilities of several “male” words are compared with the sum of an analogous set of “female” words.
Within the LLM paradigm, several clearly task-based benchmarks and evaluations have been proposed. For example, Wan et al. (2023) develops a technique for measuring bias in text generation of letters of recommendation. De-Arteaga et al. (2019) provides a benchmark for bias in classification and prediction of gender in professional biographies. As discussed, Parrish et al. (2022) developed a widely popular benchmark for bias and stereotyping in question answering. And, Zhao et al. (2018) compiled the WinoBias benchmark, a dataset measuring gender bias in coreference resolution. Each of these have remained firmly in the extrinsic camp.
@@ 2.3 Fine-tuned models
Finally, as fine-tuning of models became more commonplace, the divide between intrinsic and extrinsic has, by some, come to be defined by whether a task is performed before fine-tuning or after fine-tuning. Ladhak et al. (2023) studied the relationship between upstream (“intrinsic”) and downstream (“extrinsic”) metrics in versions of BART (Lewis et al., 2020) and PEGASUS (Zhang et al., 2020) that were fine-tuned for text summarization. The upstream metric was based on the pre-trained base model’s ability to correctly state a person’s nationality when prompted with <name> is a citizen of. The downstream task was based on perturbed descriptions of individuals, which replaced the name of a person of one nationality with the name of a person of another nationality.
In the upstream-downstream paradigm used in these papers, even benchmarks that were clearly extrinsic before, such as BBQ, can now be considered intrinsic if the evaluations are run prior to fine-tuning. However, it may be difficult to apply this paradigm to modern LLM use cases in which fine-tuning is no longer as common; instead, models are often used either out-of-the-box, with a system prompt (i.e., a prompt prepended to all user prompts), or with in-context learning (i.e., task-specific examples and explanations in the user prompt).
2.4 Beyond the Intrinsic/Extrinsic divide
Despite the historical emphasis on the distinction between intrinsic and extrinsic evaluations, there are good arguments in favor of developing additional ways to describe and categorize AI evaluations.
The first is that with the advent of newer generations of language models, the distinction between intrinsic and extrinsic evaluations is no longer crisp and well-defined, even if we set aside the social dynamic of definitions shifting over time. For example, if we consider evaluations based on log probabilities of words in sentences to be intrinsic, there are ways to formulate these tasks so that they could be considered an extrinsic text generation task instead. For example, consider the Gender Sensitivity English evaluation, which calculates the next-word probability of several “male” and “female” words in a simple sentence about an occupation. One could directly extract the next word probabilities as defined in the original test, or one could create many different sentence completions. By counting the number of times that any of the “male” and “female” words appeared in the generated outputs, experiments would asymptotically arrive at the same information as the extracted token probabilities. In other words, one could arrive at the same metric by empirically counting the number of times the next word in the sentence completion was in the “male” and “female” categories after many trials. Would the Gender Sensitivity English evaluation then be extrinsic instead of intrinsic? Likewise, extrinsic evaluations–at least those that are dependent on the appearance of a single token in text generation–could theoretically be generated intrinsically by examining the probabilities directly. Extrinsic evaluations based on the generation of more than one token could be, and in fact are, based on repeated conditional probabilities of token generation, either taken individually or aggregated with techniques such as beam search and top-K sampling. The line between intrinsic and extrinsic, in many cases, seems more dependent on the details of implementation rather than the content of the evaluation itself.
Second, even among “extrinsic” evaluations, there is a wide degree of variability in how realistic the use case representing the task actually is. To take the BBQ benchmark as an example, while question answering is certainly a frequent model task, we can imagine few cases in which a user— not just a researcher— would instruct the model to guess whether, for instance, a generic “girl” or “boy” is better at math with only a tiny amount of information about them. That is, even though the popular BBQ benchmark and other similar bias benchmarks are technically extrinsic (the implied gold standard in studies of the intrinsic-extrinsic relationship), they are based on isolated and abstract snippets of text that are decontextualized from realistic use and disconnected from real harm. They are cleverly designed to elicit stereotypical or “biased” output in abstract settings. Because the evaluations are so decontextualized, they seemingly provide proxies for generalized model behavior across contexts or evidence of an underlying model “belief”.
end of this spectrum are one-off bias evaluations based on realistic use that are directly connected to how the manifestation of bias in that context translates to actual, real world harm. One recent study that most closely meets the RUTEd ideal is Wan et al. (2023), which studies bias in generated letters of recommendation.
A substantial body of previous work has thoroughly explored the relationship between intrinsic and extrinsic evaluations (Goldfarb-Tarrant et al., 2020; Cao et al., 2022; Kaneko et al., 2022; Delobelle et al., 2022; Jin et al., 2021; Ladhak et al., 2023) and largely found that intrinsic metrics offer little utility for evaluating bias in downstream use. Our work follows in this tradition of comparing whether different types of evaluations provide information about one another, focusing instead on the RUTEd axis.
@@ 3 Evaluations
For the present work, we adapt three de-contextualized trick test evaluations from the literature. We compare the results of these evaluations with three RUTEd evaluations of long-form content generation tasks. We focus exclusively on gender-occupation bias for binary gender. While there are many other types of bias that could arise and we acknowledge the limitations of binary gender, we choose this paradigm for the sake of internal consistency across evaluations and consistency with prior work. This paradigm is so common that Weidinger et al. (2023) noted that a full 17% of “bias” benchmarks in their survey of evaluations specifically pertain to gender-occupation bias.
3.1 Trick Test Tasks
For our first task, we follow recent discussions of bias in state-of-the-art LLMs (Kapoor, 2023) that utilize the WinoBias sentences referring to occupations and gendered pronouns, such as the sentence, “The physician hired the secretary because he was overwhelmed with clients.” For each sentence in the benchmark corpus, we input the prompt
{sentence}. Who does {pronoun} refer to? Answer:
and extract the next-word probabilities for both possible occupations (“physician” and “secretary”). We normalize these probabilities such that, for each test sentence, they sum to one and retain the probability of the correct answer (“physician”). For sentence $s$, we call this probability $p^{s}_{s}$.
We also test the alternative sentence, for which the correct answer is inconsistent with prevailing gender stereotypes (“The physician hired the secretary because she was overwhelmed with clients”). We calculate the analogous probability that “she” refers to the physician. For sentence $s$, we call this value $p^{a}_{s}$. Here, $a$ refers to the fact that this sentence is anti-stereotypical.
This benchmark contains $3,160$ sentences for evaluation. The WinoBias benchmark also contains a list of 20 female-majority occupations and 20 male-majority occupations that were used to create its test sentences. We use this list of occupations in the other tasks for consistency.
Our second trick test is based on the neutrality test in the Gender Sensitivity-English task (Srivastava et al., 2022). For this task, for each occupation in the list of occupations, we input the prompt:
@@ 3.1 Trick Test Tasks
and calculate the next-word probabilities of a set of male words and a set of female words. The male words are [’male’, ’he’, ’his’, ’him’, ’himself’, ’himselves’, ’masculine’, ’masculinity’, ’masculinities’, ’boy’, ’boys’], and the female words are [’female’, ’she’, ’hers’, ’her’, ’herself’, ’herselves’, ’feminine’, ’femininity’, ’femininities’, ’girl’, ’girls’]. For example, this test calculates the probability under the model that the next word in the sentence “accountant is” is “his.” For each occupation, the probabilities of all of the male words are summed, and the probabilities of the female words are summed to arrive at the probability that the next word is in the male or female set, respectively. These probabilities are then normalized such that the probability of a male word and the probability of a female word sum to one. For the sentence pertaining to occupation $o$, we call the normalized probability that the next word is in the male set $p^{m}_{o}$ and the normalized probability that the next word in the sentence is in the female set $p^{f}_{o}$.
Although the original benchmark contained its own set of occupations, for consistency across tasks, we use the list of occupations provided in the WinoBias dataset.
3.2 RUTEd Tasks
We propose three long-form text generation that we believe are more connected to typical LLM use and for which gender-occupation bias in this context could result in a tangible problem. Example outputs for each task are provided in the appendix.
3.2.1 Bedtime Stories
Despite their clearly widespread usage, public data on the particular use cases most common for LLMs is limited. Still, one provocative example that became frequently discussed in 2023 is generating bedtime stories for small children (BedtimeStory.ai, 2023; Kobie, 2023; McGuinness, 2023; OpenAI, 2023; Srivastava, 2023). Bedtime stories are a daily interaction for many parents, and storytelling has the ability to spark a child’s imagination and shape what they think of as possible. For this reason, we believe that reinforcing stereotypes—in our case, between gender and occupation—may directly lead to real-world harm. It may might subtly influence children’s beliefs about the types of occupations available to them.

### v2

# Bias in Language Models: Beyond Trick Tests and Towards RUTEd Evaluation

@@ Abstract
Standard benchmarks of bias and fairness in large language models (LLMs) measure the association between social attributes implied in user prompts and short LLM responses. In the commonly studied domain of gender-occupation bias, we test whether these benchmarks are robust to lengthening the LLM responses as a measure of Realistic Use and Tangible Effects (i.e., RUTEd evaluations). From the current literature, we adapt three standard bias metrics (neutrality, skew, and stereotype), and we develop analogous RUTEd evaluations from three contexts of real-world use: children’s bedtime stories, user personas, and English language learning exercises. We find that standard bias metrics have no significant correlation with the more realistic bias metrics. For example, selecting the least biased model based on the standard “trick tests” coincides with selecting the least biased model as measured in more realistic use no more than random chance. We suggest that there is not yet evidence to justify standard benchmarks as reliable proxies of real-world biases, and we encourage further development of context-specific RUTEd evaluations.
Bias in Language Models:
Beyond Trick Tests and Towards RUTEd Evaluation
Kristian Lum1, Jacy Reese Anthis1,2, Kevin Robinson3, Chirag Nagpal3, Alexander D’Amour4 1University of Chicago, 2Stanford University, 3Google Research, 4Google DeepMind
@@ 1 Introduction
As large language models (LLMs) are increasingly used in everyday life, numerous concerns have been raised about the ethical impacts on users and society at large. From these concerns have sprung a number of benchmarks to assess bias and fairness in LLMs Anthis et al. (2024); Gallegos et al. (2023). Standard bias benchmarks are built on testing the correlation between sensitive attributes and other social attributes, typically gender (e.g., gendered pronouns) and occupation (e.g., manager, nurse). While the underlying social associations are complex and highly context-dependent, the benchmark inputs and outputs are typically brief, such as the probability of completing the phrase, “Nurse is,” with either a word associated with men or a word associated with women.
These benchmarks have been criticized for unstated assumptions, a lack of motivation, and conceptual issues (Blodgett et al., 2020, 2021). Yet, such benchmarks are still the predominant form of bias assessment for LLMs. For example, the Flan-PaLM models developed by Google and the Claude models developed by Anthropic were both tested with one such benchmark, the Bias Benchmark for Question Answering (BBQ), and a reduction in BBQ score was described as an improvement in bias from past model versions (Anthropic, 2023; Google, 2022).
We have very little empirical understanding of how well such bias benchmarks predict real-world bias and harm, particularly in context-specific use cases of text generation. Previous work has divided bias metrics primarily between “intrinsic” metrics—more associated with the initial representations and behavior of models—and “extrinsic” metrics—more associated with downstream model behavior (Goldfarb-Tarrant et al., 2020; Cao et al., 2022; Kaneko et al., 2022; Delobelle et al., 2022; Jin et al., 2021; Ladhak et al., 2023). This work has argued that intrinsic metrics offer little utility for evaluating bias in downstream use, but as we will evidence, this distinction has limited utility in LLM evaluation because there is little evidence that even extrinsic metrics predict more realistic task performance.
We argue that standard benchmarks constitute “trick tests”: decontextualized evaluations based on contrived scenarios designed to elicit a simplified correlation between model output and a sensitive attribute rather than as best estimates of the real-world effects of model use. We contrast these tests with novel evaluations that are grounded, at least to some extent, in Realistic Use and Tangible Effects, or RUTEd (“rooted”) evaluations. The need for RUTEd evaluations echoes calls for sociotechnical evaluations of ML systems, beyond the current focus on “a small space of the potential harm” (Weidinger et al., 2023). We conduct this study in the context of gender-occupation bias, the most common association tested in bias benchmarks (Weidinger et al., 2023), and this association allows us to sidestep much of the subjectivity and debates around other social contexts, such as race and socioeconomic status (Blodgett et al., 2021).
Among the nine LLMs that we tested, if one used standard benchmarks to guess which candidate model is the least biased in the long-form text evaluations, they would do no better than random chance. Further, bias evaluations in each context were largely uncorrelated with each other, suggesting that bias measured in one context may not reliably generalize to other contexts. Rather, addressing LLM bias may require bespoke evaluations based on particular uses and affected populations. More research is needed to understand, measure, and address LLM bias—especially work that measures not just realistic use, but tangible effects, by conducting human subjects research.
In this paper, we use the following terminology. An evaluation is the application of a metric to a particular task. A task is a combination of a prompt and the dataset on which the model is tasked with implementing that prompt. A metric is a formula that summarizes the model’s performance at that task. When an evaluation becomes standardized (e.g., compared to other evaluations, published in a peer-reviewed venue), it is often described as a benchmark. In Section 2, we review the intrinsic-extrinsic metric distinction and motivate a more grounded conceptualization. In Section 3, we develop the RUTEd framework in contrast with standard benchmarks, and we present our results in Section 4 before concluding and outlining limitations.
2 Intrinsic and extrinsic bias evaluations
The meaning and measurement of bias has long been critiqued and contested in the NLP literature. Blodgett et al. (2020) reviewed use of the term “bias,” finding that researchers use a wide range of normative motivations—often only briefly or vaguely specified—including stereotyping, questionable correlations between model behavior and language features, allocational harms (e.g., the distribution of jobs or financial resources), and a nebulous category of other representational harms (e.g., system performance, misrepresentation, denigration). Likewise, Blodgett et al. (2021) argued that common benchmark datasets have a number of pitfalls, such as conflating race, culture, and nationality as well as logical and grammatical issues.
While debates about the fundamental definitions of bias and fairness are beyond the scope of this work, our work builds on the distinction between intrinsic and extrinsic bias metrics. As originally defined by Goldfarb-Tarrant et al. (2020), intrinsic metrics measure properties inherent to the model, and extrinsic bias evaluations measure the biases relative to a specified task. However, the usage of these terms has changed significantly over time, suggesting the need for new conceptualizations.
2.1 Static word embeddings
@@
Popular intrinsic bias metrics of this sort include the Word Embedding Association Test (WEAT) benchmark Caliskan et al. (2017)) and the similar approach of Bolukbasi et al. (2016). Both aggregate cosine similarity measures between words associated with different identity groups (e.g., “he,” “she”) with words in a domain of interest (e.g., occupations). In the paper that introduced the intrinsic-extrinsic dichotomy (Goldfarb-Tarrant et al., 2020), the intrinsic metric of WEAT (in both English and Spanish) was contrasted with extrinsic metrics based on models that used those embeddings for the tasks of coreference resolution and hate speech detection.
@@ 2.2 LLMs
As the dominant NLP paradigm shifted towards LLMs, so did what is considered “intrinsic.” In contrast to static word embedding models, LLMs contain dynamic embeddings that change with context. To evaluate bias in this paradigm, Guo and Caliskan (2021) developed an extension of WEAT, the Contextualized Embedding Association Test (CEAT). Another paper on the intrinsic-extrinsic connection, Cao et al. (2022), adapted to this shifting paradigm with numerous experiments on 19 models, primarily variants of BERT and GPT-2. In this study, they considered CEAT and two other benchmarks–StereoSet (Nadeem et al., 2020) and ILPS (Kurita et al., 2019)–as “intrinsic metrics,” even though they are not based on the embedding space itself but on the log probabilities of words in text that can evoke stereotypes. These probabilities constitute task performance in the sense that they reflect the next-word predictions of a non-zero temperature LLM over many trials.
Several task-based evaluations have been developed, which go beyond single-word outputs. For example, Wan et al. (2023) develops a technique for measuring bias in generated letters of recommendation. De-Arteaga et al. (2019) provides a benchmark for bias in classification and prediction of gender in occupational biographies. As discussed, Parrish et al. (2022) developed a widely popular benchmark for bias and stereotyping in question answering. And, Zhao et al. (2018) compiled the WinoBias benchmark, a dataset measuring gender bias in coreference resolution.
@@ 2.3 Fine-tuned models
Finally, as fine-tuning of models became more commonplace, the divide between intrinsic and extrinsic has, by some, come to be defined by whether a task is performed before or after fine-tuning. Ladhak et al. (2023) studied the relationship between upstream (“intrinsic”) and downstream (“extrinsic”) metrics in versions of BART (Lewis et al., 2020) and PEGASUS (Zhang et al., 2020) that were fine-tuned for text summarization. The upstream metric was based on the pre-trained base model’s ability to correctly state a person’s nationality when prompted with <name> is a citizen of. The downstream task was based on perturbed descriptions of individuals, which replaced the name of a person of one nationality with the name of a person of another nationality. The downstream metric was the hallucination rate, defined as a model incorrectly summarizing the description by stating that the person was of the original nationality rather than the one in the new description. For example, a model hallucinates if the name of a Japanese person, “Naoki Tsukahara,” is inserted into the biography of a French person that mentions they are from France, but the model states that Naoki Tsukahara is from Japan.
2.4 Beyond the intrinsic-extrinsic divide
For modern LLMs, the intrinsic-extrinsic divide may be more useful if reframed as a wide spectrum, ranging from the embedding space within a model to the most downstream use after fine-tuning and instruction-tuning (e.g., with RLHF). Still, it is difficult to firmly place evaluations on this spectrum because, as described, more intrinsic metrics (e.g., word probabilities) can be translated into apparently extrinsic metrics (e.g., text generation).
Moreover, even extrinsic evaluations usually seem unrealistic. To take the BBQ benchmark as an example, the extrinsic task of question answering—extrinsic in that it is about model behavior rather than internal representation—is a frequent LLM use, but there are few cases in which a user would instruct the model, as in BBQ, to guess whether a generic “girl” or “boy” is better at math. While extrinsic, it is nonetheless based on isolated and abstract snippets of text that have no verified connection to real-world use and harm.
Several studies have empirically explored the correlations between evaluations at different points along the intrinsic-extrinsic spectrum (Goldfarb-Tarrant et al., 2020; Cao et al., 2022; Kaneko et al., 2022; Delobelle et al., 2022; Jin et al., 2021; Ladhak et al., 2023). These studies have largely found that more intrinsic or upstream metrics offer little utility for evaluating extrinsic or more downstream bias. Our work builds upon this, but our findings suggest a stronger claim: Even extrinsic metrics typically do not reflect realistic use or relate clearly to tangible effects, and extrinsic metrics fail to predict the extent of LLM bias in even relatively simple use cases of text generation.
Nonetheless, standard benchmarks have been and will continue to be useful. One reason is that more complex prompt-based tests require more “metalinguistic judgment”—such that they do not always reveal LLM capabilities that would show up in rigorous testing of their internal representations (Hu and Levy, 2023).
@@ 3 Evaluations
We adapt a standard benchmark from the literature,the Gender Sensitivity task from the popular BIG-bench testing suite (Srivastava et al., 2022), and compare it with three long-form text generation tasks based on realistic use cases. Each evaluation has three derived metrics, and we test nine LLMs. We focus exclusively on the context of gender-occupation bias associated with male and female genders. While there are many other types of bias that could arise, and we acknowledge the limitations of the binary gender framework, we chose this paradigm for internal consistency across evaluations and consistency with prior work. Gender-occupation is the most prevalent paradigm in the literature on bias in NLP with approximately 17% of bias evaluations using it according to a recent literature review (Weidinger et al., 2023).
3.1 Decontextualized standard benchmark tasks
Our decontextualized evaluation was based on the neutrality test in the BIG-bench Gender Sensitivity-English task (Srivastava et al., 2022). For comparability with other evaluations, we do not use the occupations provided for this task but instead the 20 female-majority occupations and 20 male-majority occupations from WinoBias (Zhao et al., 2018). We run the test by inputting the prompt:
@@ 3.1 Decontextualized standard benchmark tasks
and calculating the next-word probabilities of each word in a set of male words and a set of female words. The male words are [‘male’, ‘he’, ‘his’, ‘him’, ‘himself’, ‘himselves’, ‘masculine’, ‘masculinity’, ‘masculinities’, ‘boy’, ‘boys’], and the female words are [‘female’, ‘she’, ‘hers’, ‘her’, ‘herself’, ‘herselves’, ‘feminine’, ‘femininity’, ‘femininities’, ‘girl’, ‘girls’]. For example, we calculate the probability that the next word in the sentence “accountant is” is “his.” For each occupation, the probabilities of male and female words are each summed and normalized. We describe the metrics calculated with these probabilities in Section 3.3. For the sentence pertaining to occupation $o$, we call the normalized probability that the next word is in the male set $p^{m}_{o}$ and the normalized probability that the next word in the sentence is in the female set $p^{f}_{o}$.
3.2 RUTEd evaluation tasks
To test whether standard benchmarks measure a stable feature of LLMs that can be directly translated to real-world contexts, we drew on media discussions to identify common use cases that are suitable for evaluation and more connected to tangible effects. Note that our evaluation tasks are only intended for research, not deployment as real-world evaluations. Example outputs for each task are provided in Appendix C.
Bedtime stories
Public data on LLM use is limited, but one example that is frequently discussed in the media is generating bedtime stories for small children (BedtimeStory.ai, 2023; Kobie, 2023; McGuinness, 2023; OpenAI, 2023; Srivastava, 2023). This is reportedly a frequent use case in which models perform relatively well, and bedtime stories are a daily interaction for many parents. Storytelling has the ability to spark a child’s imagination and shape what they think of as possible. For this reason, we believe that reinforcing stereotypes—particularly in the most common bias example of gender-stereotyped occupations—may directly lead to real-world harm. It might subtly influence children’s beliefs about the types of occupations available to them.